### v2

# Identifying User Goals from UI Trajectories

Abstract
Autonomous agents that interact with graphical user interfaces (GUIs) hold significant potential for enhancing user experiences. To further improve these experiences, agents need to be personalized and proactive. By effectively comprehending user intentions through their actions and interactions with GUIs, agents will be better positioned to achieve these goals. This paper introduces the task of goal identification from observed UI trajectories, aiming to infer the user’s intended task based on their GUI interactions. We propose a novel evaluation metric to assess whether two task descriptions are paraphrases within a specific UI environment. By Leveraging the inverse relation with the UI automation task, we utilized the Android-In-The-Wild and Mind2Web datasets for our experiments. Using our metric and these datasets, we conducted several experiments comparing the performance of humans and state-of-the-art models, specifically GPT-4 and Gemini-1.5 Pro. Our results show that Gemini performs better than GPT but still underperforms compared to humans, indicating significant room for improvement.
Identifying User Goals from UI Trajectories
Omri Berkovitchthanks: Equal contribution, 1  Sapir Caduri11footnotemark: 1, 1  Noam Kahlon1  Anatoly Efros1 Avi Caciularu1Ido Dagan1, 2 1Google Research  2Bar-Ilan University {berkovitchomri,sapir,kahlonn,talef,avica,idodagan}@google.com
1 Introduction
Autonomous agents that interact with GUIs to complete tasks for users have drawn significant interest Hong et al. (2023); Gur et al. (2023); Yang et al. (2023b). These agents interpret natural language task descriptions provided by a user and iteratively interact with GUIs to complete the desired task. Enhancing an agent’s ability to understand user actions and infer goals within the GUI environment has the potential to significantly increase its utility for users by providing more personalized and effective assistance Li et al. (2024).
Consider the scenario in Figure 1 where a user books flight tickets online for a vacation. An ideal agent would first observe and understand these actions, then proactively suggest booking a hotel for the same dates, and make the vacation dates visible in the calendar. This kind of assistance requires the agent to understand the underlying goals driving user interactions. Additionally, by identifying user goals, we can augment the agent’s memory, making it more personal and useful for future tasks. We frame this challenge as the inverse of the known UI automation task: identifying the hidden user goal from its observed UI interactions Li et al. (2023); Wen et al. (2023). While intent identification has been explored in other domains Hu et al. (2009); Gupta et al. (2014), our work is the first, to the best of our knowledge, to focus specifically on identifying user goals from UI interactions.
Figure 1: An example of a user performing a flight booking task. The agent first observes the UI interactions, comprehends the task’s essence, and then offers help with related tasks, like booking a hotel and blocking calendar dates. We focus on the first part, comprehending the task by observing the UI interactions.
Identifying user goals from UI interactions is challenging. Although ensuring a correct goal is achievable through the observed actions is essential, this alone is not enough. The fact that multiple tasks can often lead to the same actions making the task ambiguous, further highlighting the complexity of the task and the need for robust methods. Additionally, the involvement of multiple modalities, such as visual and text, increases complexity.
In this paper, we explore the task of goal identification from observed UI trajectories. In particular, we introduce a robust manual and automatic evaluation protocol. Our proposed metric primarily assesses whether two task descriptions are paraphrases within the UI context while also enabling us to quantify the ambiguity. We conducted extensive experiments in both web and Android UIs, leveraging existing UI automation datasets by swapping the roles of input and output. We compared the performance of humans and state-of-the-art multi-modal models on this task and found that the models significantly underperform, indicating substantial room for improvement. Common errors and potential mitigations are highlighted, paving the way for future research.
Overall our main contributions are as follows: (1) presenting the first study on the task of goal identification from UI trajectories, including task formulation and definitions; (2) introducing a novel manual and automatic evaluation methodology and (3) assessing the performance of both humans and models on this task, providing a comprehensive analysis of our findings.
2 Task Definition
Given an observed UI trajectory performed by a user with the intention to complete a certain task, our goal is to recover the user’s original intent from the observable trajectory. As mentioned in § 1, this intent identification setting is effectively the inverse problem of the known UI Automation task. We therefore adopt their input and output definitions, swapping their roles, which enables the use of UI automation datasets for the intent identification task as well. Accordingly, our input is a UI trajectory - a sequence of steps performed by the user. Each step consists of a snapshot of the UI environment at that moment, along with the corresponding action the user took at that step. From this trajectory, our goal is to generate a natural language description that accurately captures the user’s intended task. Within the scope of this paper, we address the core setting of the intent identification task and therefore assume that the observed UI trajectory indeed successfully fulfills the underlying user intent.
The intent identification task, similar to other text generation tasks like summarization where multiple valid outputs can exist, is inherently ambiguous, mostly because the same trajectory may fulfill multiple intents. For example, when observing a Sushi restaurant booking, the user might have asked for that specific Sushi place, or more broadly for some restaurant in that area, or of that cuisine type. When a model identifies intent from trajectory, we expect it to predict the most likely one.
3 Evaluation Methodology
This section outlines our evaluation methodology. Given an input UI trajectory and a corresponding gold task description, we assess whether a predicted task description matches the gold reference. We start with necessary definitions (Section 3.1), followed by a human evaluation protocol (Section 3.2) and a language model-based automatic evaluation (Section 3.3).
3.1 Definitions
Task Fulfillment by a Trajectory
As mentioned in § 2, we assume an observed UI trajectory fulfills the underlying user’s intended tasks. Therefore, a predicted task description not fulfilled by the input trajectory is erroneous and does not match the gold task description.
Inspired by the taxonomy proposed by Zhou et al. (2023), we differentiate between fulfillment for information-seeking intents and transactional intents (e.g. purchasing an item, booking a flight or changing settings). For the latter, fulfillment is achieved upon successfully completing the specific requirement outlined in the task. For information-seeking intents, fulfillment is achieved when the trajectory provides the necessary information sought in the user intent. We note that a fulfilling trajectory may provide some additional information beyond the intent, if such additional information is inherently bundled in the UI environment together with the sought information (e.g. some UI environments may display review scores along with other related information).
Satisfaction Relation between Tasks
We next aim to specify a matching criterion between two task descriptions, specifically a predicted one and the corresponding gold-reference. Given two task descriptions A and B and a UI environment, we say that A satisfies B in that environment if every reasonable user trajectory that fulfills A would also fulfill B. In essence, this means that completing task A necessarily results in completing task B, making B a more general task than A in that UI environment. For instance, the task “Purchase the earliest train ticket to Edinburgh” satisfies the task “Purchase a ticket to Edinburgh” but not vice versa.
Building on these definitions, we consider a predicted task description to successfully match the gold description if the two mutually satisfy each other, and partial match when only one satisfies the other. Essentially, matching tasks can be considered as paraphrases of the same intent within the context of the UI environment. For instance, the tasks “Find a large dining table” and “Find a dining table for 10-12 people” match each other if the UI environment considers 10-12 people as large. We highlight the relation between match and ambiguous trajectory, which fulfills multiple task descriptions that do not match each other.
3.2 Human Evaluation Protocol
As with many text generation tasks, human evaluation is essential due to the limited reliability of automatic evaluation metrics. In our case, the human evaluator observes the gold and predicted task descriptions, along with the corresponding trajectory, and assesses (1) whether the trajectory fulfills the predicted task, and (2) whether each of the predicted and gold task descriptions satisfies the other.
To verify the quality of our suggested metric, we sampled randomly 50 instances from Mind2Web dataset (see § 4.1) and generated task descriptions using two baseline models (see § 4.2) evaluated in our experiments. We measured the pairwise inter-annotator agreement between three of the authors yielding average Cohen’s Kappa of 0.79 for fulfillment judgement and 0.77 for satisfaction judgement. These agreement levels are in the higher range of the “Substantial agreement” level, thus we proceeded to manually evaluate each baseline model in our experiments § 5 by a single annotator.
3.3 Automatic Evaluation Metric
We propose utilizing a Large Multimodal Model (LMM) as an automatic evaluator for the satisfaction criteria. Recent advancements in LMMs, such as those demonstrated in He et al. (2024) and Pan et al. (2024), have shown promising results in employing GPT-4 Achiam et al. (2023) to assess task completion by autonomous agents. Building on this, we leverage the latest GPT-4o model as the automatic evaluator to determine whether two task descriptions, are mutually satisfied in the context of the trajectory. The specific prompt used is detailed in Appendix D. Our analysis indicates that the proposed automatic evaluation generally agrees with the human annotators, with an F1 score of 0.75, and can be a sufficient substitute for manual evaluation, enabling quick development cycles.
Dataset Non Partial  Match ($\uparrow$)
 Match  Match
Mind2Web H   vs. G 0.07 0.13 0.80
  H1 vs. H2 0.09 0.10 0.81
AitW H   vs. G 0.12 0.32 0.56
  H1 vs. H2 0.05 0.20 0.75
Table 1: Manual evaluation of human generated task descriptions, at the different levels of match criteria (determined by the satisfaction relation) between annotators (Hi) and gold task descriptions (G), as well as among annotators. (H) represents average scores.
4 Datasets and Baseline Models
Manual Eval Automatic Eval
Dataset Model Fulfillment Non Match Partial Match  Match ($\uparrow$) Non Match Partial Match  Match ($\uparrow$)
Mind2Web GPT-4 0.86 0.14 0.42 0.44 0.18 0.40 0.42
Gemini-1.5 0.87 0.20 0.22 0.58 0.18 0.28 0.54
AitW GPT-4 0.65 0.50 0.20 0.30 0.62 0.19 0.19
Gemini-1.5 0.89 0.19 0.31 0.50 0.36 0.29 0.35
Table 2: Manual and automatic evaluation scores of fulfillment relation and the different levels of match criteria (determined by the satisfaction relation), between model predictions and gold task descriptions.
4.1 Evaluated Datasets
Given that our task is the inverse of UI automation task, leveraging the datasets created for UI automation is a natural choice. In these datasets, humans perform actions within a UI environment, based on a given task description.
For our experiments, we explore two UI environments: web and Android. We utilize the Mind2Web dataset Deng et al. (2023) for the web environment, as it is the most widely used benchmark for autonomous web agents. The dataset was created by curating diverse tasks across popular websites, with annotators performing a series of actions to complete the goal. For Android, we used the Android in the Wild (AitW) Rawles et al. (2023) dataset, notable for its diverse task descriptions spanning various domains. Further details in Appendix A.
4.2 Models
Given the multimodal nature of UI trajectories, our models must be adept at interpreting both text and images. We selected two state-of-the-art LMMs, Gemini 1.5 Pro Reid et al. (2024) and GPT-4 Achiam et al. (2023). These models are at the forefront of handling combined text and image inputs and offer a sufficiently large context window, which is necessary for our experiments.
Our experiments focused on in-context learning, employing both zero-shot and few-shot prompting. The model was guided through a Chain-of-Thought Wei et al. (2022) process, to analyze the trajectory in a step-by-step manner (details in Appendix D).
In web, inspired by SeeAct Zheng et al. (2024) we drew a red bounding box around the element that the user interacted with to guide the model’s attention, as well as to break ambiguity where textual action descriptions were not sufficient. Additionally, we truncated lengthy web page images based on the bounding box position to further focus the model’s attention (details in Appendix B). For Android, no special modifications were necessary, as actions were overlaid on the screenshots and the screenshots were limited to the visible area.
5 Experiments
5.1 Human Performance Evaluation
To asses task difficulty and establish a baseline for models, two NLP practitioners, unfamiliar with the datasets, independently composed 50 task descriptions from Mind2Web, and 50 from AitW. An annotator then evaluated [^1] During the evaluation it was observed that human-generated tasks were inherently fulfilled by the trajectory. Therefore, we do not report this score. these descriptions as explained in § 3.2.
The results, summarized in Table 1, reveal more matches between the human annotators and the gold task descriptions in Mind2Web compared to AitW. Upon analysis of the AitW disagreements, we found that the gold descriptions were often more specific than those provided by human annotators. In some cases, the gold tasks were already fulfilled by the trajectory, resulting in no clear interactions that indicate the user goal. For example, if the task was “Turn WiFi on” and the WiFi was already on, the annotators inferred a more general task, such as “Show WiFi settings”. Further analysis of non-matching records across both datasets revealed that ambiguous task descriptions contributed to disagreements between raters. Detailed analysis and examples can be found in Appendix C.
5.2 Model Performance Evaluation
Model evaluation included manual assessment of 100 predicted task descriptions and automated evaluation of 1,000 predictions per model and dataset (details provided in Appendix B).
The results, outlined in Table 2, reveal that Gemini outperforms GPT-4 on both datasets, achieving higher match scores but still falling short of human performance. In AitW, GPT-4 often generated too abstract goals such as “Explore the main page of a shopping app” resulting in higher non-match scores. In contrast, Gemini tended to be more specific and heavily influenced by the content on the screen. Both models, in certain instances, misidentified the actual user intent (e.g. “Watch top rated movie trailer” vs. “Watch The Dark Knight trailer”) and exhibit limitations in visual screen understanding, leading to missing details, incorporating irrelevant information and hallucinations of non-existent information. We provide a detailed error analysis and further discussion in Appendix C.
Experiments with few-shot prompting did not yield improved results. We hypothesize that the large context size, resulting from including multiple trajectories with several images each in the prompt, were too complex models to leverage.
6 Conclusion and Future Work
In this paper, we introduced a novel task for UI-related agents by identifying user goals from UI trajectories and proposed a reliable evaluation methodology. Our experiments on Android and web datasets reveal a significant gap between humans and state-of-the-art models, with a detailed analysis provided for this discrepancy.
Future work may include fine-tuning experiments and enhancing visual understanding for UI environments. We propose testing our task by evaluating how well it improves downstream tasks like agent personalization and completions. Additionally, expanding the scope to other GUIs like iOS and Windows would broaden the applicability and impact of our findings.
Limitations
Our study is subject to several limitations due to the nature of the datasets used. First, both datasets primarily include English-language websites accessed in the U.S., thereby limiting the study to English-language interactions only. Second, in real-world scenarios, we believe that (1) user trajectories may be interleaved between multiple tasks as users adjust their objectives in real time or are interrupted by other tasks, (2) users might have more ambiguous goals that evolve during their interaction with the user interface and (3) users might be less proficient with computers or phones, leading to noisier trajectories that are more challenging to identify and interpret intent from. Moreover, in the Mind2Web dataset, all tasks were limited to interactions within the same website, not encompassing multi-website tasks. This constraint may not fully represent the complexity of real-world web usage.
Lastly, this study focuses solely on Android and web environments. These environments might exhibit different task distributions compared to other user interface environments such as iOS and Windows, potentially limiting our findings for Android and web environments.
Ethical Considerations
The development of autonomous agents, while holding great potential for innovation, raises important ethical considerations. Our research focuses on understanding user intent from recorded UI trajectories, and it is crucial to acknowledge the potential privacy implications of tracking user activity. Ensuring the security and protection of this sensitive data is of the utmost importance. Employing techniques like on-device processing, anonymization, or other privacy-preserving methods can help mitigate risks and ensure user data remains protected. It is essential for researchers and developers to proactively address these concerns to foster trust and responsible innovation in the field of autonomous agents.
Acknowledgements
We wish to thank Guy Rom, Xiang Deng, and Eliya Nachmani for their valuable feedback on our research. We also would like to express our gratitude to Noam Etzion-Rosenberg and Nir Shemy for their work in data annotation.
Appendix A Datasets Overview
Figure 2: An instance from Mind2Web, representing a full trajectory accomplishing the task description above.
Figure 3: An instance from AitW, representing a full trajectory accomplishing the task "Set an alarm". The blue plus sign indicates the area on the screen where the tap occured.
Mind2Web
We aim to give a brief overview of how Mind2Web was collected, and its format in more depth. The data collection process involved selecting numerous popular websites across five top-level domains. Annotators, guided by seed tasks from GPT, proposed diverse tasks for each website. They demonstrated and recorded how to complete these tasks, resulting in a trajectory of actions and screenshots. Each action is defined by a pair consisting of a Target Element and an Operation. The Target Element is an interactable element on the current web page, such as buttons, input fields, or drop down menus. The Operation refers to the specific action to be executed on the Target Element, with Mind2Web supporting three primary operations: Click (including actions like clicking, hovering, and pressing Enter), Type (which involves entering text into input fields and requires an additional value for the text to be typed), and Select Option (which involves selecting an option from a drop down menu or similar element and also requires an additional value for the option to be selected). Notably, these actions were automatically produced during the time users were recorded completing the task, eliminating the need for additional manual labor. Figure 2 demonstrates a single instance of the data, where textual action descriptions are presented below the corresponding screenshot.
Android-In-The-Wild
The Android in the Wild (AitW) dataset stands out due to its extensive variety of tasks, covering 4 domains: Google-Apps, Install, Web-Shopping, and General, as well as a single-step domain excluded from this paper’s analysis. The dataset consists of a substantial collection of high-level instructions, trajectories of varying lengths, and a notable variety of apps and websites. A key strength of the AitW dataset lies in the diversity of task instructions, which provided in-depth insights into the complexities of understanding user goals across different domains.
The high-level goal instructions were sourced from various sources, including humans (both crowd-sourced and the authors themselves), LLM-generated prompts, and technical documentation such as PixelHelp Li et al. (2020). The creation process involved human annotators performing tasks on Android emulators, with their gestures being recorded.
The AitW dataset contains 700k+ episodes comprising 30k unique instructions. These episodes were recorded on mobile devices running four different versions of Android. Each episode contains natural language instructions and observation-action pairs. The observations are screenshots, while the actions are one of three types: tap, drag, or typing. Gesture actions are represented as taps and drags at specific <x,y> coordinates on the screen.
In our utilization of this data, we presented the models with a series of screenshots with the actions drawn on them, as shown in Figure 3.
Appendix B Experiments
Models Configuration
For task description prediction, we used the most recent version of Gemini, namely Gemini 1.5 Pro, updated in Vertex AI [^2] https://cloud.google.com/vertex-ai as of May 2024, with a sampling temperature of 1.0, and GPT-4-Turbo [^3] https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 (version gpt-4-turbo-2024-04-09) with a sampling temperature of 0.6. For automatic evaluation, we utilized the latest release from the GPT family, known as GPT-4o.
Data
For manual evaluation, we sampled 100 random data points from the test sets of both AitW and Mind2Web. For automatic evaluation, we used the entire Mind2Web test set, consisting of 1013 examples, and randomly sampled 250 examples from each of the four domains in AitW, resulting in a total of 1000 examples (as described in Appendix A).
Web
We encountered a technical challenge with the Mind2Web data due to the nature of its image captures. Unlike standard viewport captures, which represent the visible area of a web page on a typical screen, the images in Mind2Web had a median height of 4200 pixels, significantly exceeding typical web page dimensions, with 20 percent of the images exceeding 7000 pixels in height. Initial tests showed that these oversized images introduced noise and negatively affected model performance.
To address this, we implemented a heuristic truncation method, reducing image height while ensuring the interaction element remained visible within the truncated image. This was achieved by utilizing the bounding box metadata provided by Mind2Web. Similar to the approach taken by Zheng et al. (2024) and introduced by Yang et al. (2023a), we drew the red bounding box around the interaction element to guide the model’s attention. Additionally, adding the bounding box helped resolve ambiguities at times that the textual action description is not sufficient. For example, sometimes action descriptions are simply empty, and does not contain any description about the element itself. While in other cases, action descriptions exist, but the textual description matches multiple element descriptions and thus results in ambiguity that is only resolved by drawing the bounding box. For example, a button labeled "Add to Cart" is typically associated with each item in a web shopping list. Without the bounding box, it is impossible to determine which specific button was clicked.
Android
In our efforts to replicate experiments from web, we encountered challenges due to differences in data format. Unlike the web, AitW does not provide textual information associated with clicked elements. Instead, it offers x,y coordinates representing the center of the tapped area. This distinction in data structure made it impractical to conduct experiments involving both actions written alongside screenshots.
To address these challenges, we utilized the dataset’s utilities to overlay actions on top of the screenshots, as well as bounding box annotations for post-process detected UI elements. However, we found that the added element annotation marks often confused the model added noise and critical information on the screen. As a result, we proceeded with experiments using a sequence of screenshots that had the actions (tap, drag, and type) drawn over them. AitW’s provided visualization tools to draw the actions, also included labels of special actions such as the back button, home button, and enter. As well as a special "status" action: either Task Complete or Task Impossible. However, we found that the label Task Complete often confused the models. Despite our efforts to make the models ignore it, we eventually abandoned this specific annotation as it is a technicality of data representation.
Appendix C Error Analysis
Figure 4: An illustration in which the user chose a specific car primarily for its 12-inch feature, but since it was also the cheapest, annotators incorrectly assumed cost was the deciding factor.
In this section, we aim to provide a more detailed error analysis with respect to model and humans.
Web
In our error analysis of 30 mismatched task descriptions, distinct patterns emerged between GPT and Gemini models. GPT frequently wrote tasks as navigational procedures, with 12% of generated tasks starting with “Navigate” (out of the 1,000 predicted tasks). Additionally, over 20% of the manually inspected errors involved misinterpreting the task’s intent, often producing broad descriptions lacking crucial details. Conversely, Gemini’s errors were typically more fine-grained, often capturing the task’s essence but struggling with specific details like dates, numbers, or locations.
Both models also occasionally produced task descriptions that felt artificial, incorporating information a user would be unlikely to know beforehand due to the models’ access to the full user trajectory. For example, a task like “Read recent news about Apple stock” might be predicted as “Read the article ’X’ about Apple stock” if the model observed the user clicking on a specific article ’X’.
With respect to human annotators, we found that most disagreements between human-generated task descriptions and gold task descriptions resulted from humans making more generalized tasks. This happened because they choose the most natural or probable constraints if no action provides evidence for a less likely constraint. Sometimes, they don’t write the constraint or any other one if nothing seems probable. Figure 4 demonstrates such a case, the truck picked by the user is the only 12-inch wheel truck but also the cheapest truck among the listed options, as no prior action gave evident to the 12-inch constraint, both human generated task labeled it as “Book the cheapest truck…” while the gold task description was “Book a 12-inch wheel truck…”. Additionally, with respect to task ambiguity, each trajectory in this experiment resulted in three task descriptions, two from the human annotators and one from the gold reference. We calculated the number of trajectories where all task descriptions matched each other and those where they did not. We found that 72% of trajectories had a match among all tasks, 24% had a match between two tasks, and the rest had no matching tasks. Although not exhaustive, this highlights that most trajectories in Mind2Web are probably non-ambiguous.
Figure 5: Comparison of "Match" proportions between Gemini 1.5 Pro and GPT-4-Turbo models across different AitW domains
Android
Comparing GPT and Gemini’s performance on the Android dataset revealed notable variations across different domains. Gemini consistently surpassed GPT in all domains, primarily due to GPT’s tendency to offer abstract goals and frequent random hallucinations. Gemini’s errors varied based on the specific domain.
Both models were proficient in the “General” domain, but faced challenges in “Web Shopping” and “Google Apps”. This disparity is due to the nature of the “General” dataset, which contains search queries that explicitly reveal the user’s intent.
“Google Apps” tasks overlap with PixelHelp, primarily involving settings configuration. Some tasks are ambiguous, such as a button toggle, but the trajectory displays only one option. Additionally, specific tasks request actions that have already been completed, e.g Turn On location history, but the trajectory only shows viewing the Location History setting page, leading to confusion for models. Additionally, the models faced difficulty comprehending the correct order of the sequence of actions that occurred, frequently mistaking the final state (on or off).
On the “Web Shopping” domain, Gemini provided excessive details about specific products (“Add Razer Kraken X for Console Gaming Headset for PC/PS4/PS5/Xbox/Switch - Black/Blue to cart on Best Buy Canada”). GPT often missed the main purpose of the task and suggested abstract tasks such as: “log into an account” or “Decline the offer to protect a purchase with an insurance plan on a shopping website”.
The “Install” domain often presents ambiguous tasks in the format “open (install if not installed)” which confuses both models. Furthermore, in some cases the apps were already pre-installed which made it impossible to predict (model predicts “open”), providing only partial satisfaction in one direction.
These results indicate that further refinement and training may be needed to improve the models’ performance in specific domains.
For human generated tasks, “General” dataset presented minimal challenges for human annotators. This was attributed to their ability to effortlessly comprehend the user’s intended intent based solely on the visible search query. However, analogous to the model challenges encountered, ambiguous tasks within the “Install” dataset proved challenging for humans as well. Conversely, unlike models, humans exhibited impeccable performance in comprehending the final state of the desired setting configuration within the Google Apps domain, if the original goal was specific and not ambiguous. Shopping tasks, on the other hand, posed a distinct challenge for humans. They struggled to grasp the rationale behind selecting an item when the original task was to choose the cheapest or the first result. These findings underscore the multifaceted nature of goal task prediction and emphasize the significance of addressing specific domains.
Appendix D Instructions and Prompts
Figure 6: Instructions for human-annotators to conclude if a task is fulfilled by a trajectory, and if two task descriptions satisfying each other. These instructions were the core prompt of the automatic evaluator.
Figure 7: Goal identification annotation task instruction to measure human performance.
Figure 8: Instruction used to guide GPT and Gemini when predicting a task description given a UI trajectory. We swap AitW section and Mind2Web according to the input.
Figure 9: Model instructions for evaluating Satisfaction relation between two task descriptions (A and B), given a corresponding trajectory. Includes web/mobile data format instructions and few-shot examples for experimentation (detailed on the next page).
Figure 10: Few-Shot exemplars demonstrating the Satisfaction Relation Prompt (Figure 9). These simplified examples highlight the nuances of task satisfaction in real-world scenarios.

### v3

@@
 Identifying User Goals From UI Trajectories
-Abstract
-Autonomous agents that interact with graphical user interfaces (GUIs) hold significant potential for enhancing user experiences.
-To further improve these experiences, agents need to be personalized and proactive.
-By effectively comprehending user intentions through their actions and interactions with GUIs, agents will be better positioned to achieve these goals.
-This paper introduces the task of goal identification from observed UI trajectories, aiming to infer the user’s intended task based on their GUI interactions.
-We propose a novel evaluation metric to assess whether two task descriptions are paraphrases within a specific UI environment.
-By Leveraging the inverse relation with the UI automation task, we utilized the Android-In-The-Wild and Mind2Web datasets for our experiments.
-Using our metric and these datasets, we conducted several experiments comparing the performance of humans and state-of-the-art models, specifically GPT-4 and Gemini-1.5 Pro.
-Our results show that Gemini performs better than GPT but still underperforms compared to humans, indicating significant room for improvement.
-Identifying User Goals from UI Trajectories
-Omri Berkovitchthanks: Equal contribution, 1  Sapir Caduri11footnotemark: 1, 1  Noam Kahlon1  Anatoly Efros1 Avi Caciularu1Ido Dagan1, 2 1Google Research  2Bar-Ilan University {berkovitchomri,sapir,kahlonn,talef,avica,idodagan}@google.com
-1 Introduction
-Autonomous agents that interact with GUIs to complete tasks for users have drawn significant interest Hong et al.
-(2023); Gur et al.
-(2023); Yang et al.
-(2023b).
-These agents interpret natural language task descriptions provided by a user and iteratively interact with GUIs to complete the desired task.
-Enhancing an agent’s ability to understand user actions and infer goals within the GUI environment has the potential to significantly increase its utility for users by providing more personalized and effective assistance Li et al.
-(2024).
-Consider the scenario in Figure 1 where a user books flight tickets online for a vacation.
-An ideal agent would first observe and understand these actions, then proactively suggest booking a hotel for the same dates, and make the vacation dates visible in the calendar.
-This kind of assistance requires the agent to understand the underlying goals driving user interactions.
-Additionally, by identifying user goals, we can augment the agent’s memory, making it more personal and useful for future tasks.
-We frame this challenge as the inverse of the known UI automation task: identifying the hidden user goal from its observed UI interactions Li et al.
-(2023); Wen et al.
-(2023).
-While intent identification has been explored in other domains Hu et al.
-(2009); Gupta et al.
-(2014), our work is the first, to the best of our knowledge, to focus specifically on identifying user goals from UI interactions.
-Figure 1: An example of a user performing a flight booking task.
+Abstract.
+Identifying underlying user goals and intents has been recognized as valuable in various personalization-oriented settings, such as personalized agents, improved search responses, advertising, user analytics, and more.
+In this paper, we propose a new task—goal identification from observed UI trajectories—aiming to infer the user’s detailed intentions when performing a task within UI environments.
+To support this task, we also introduce a novel evaluation methodology designed to assess whether two intent descriptions can be considered paraphrases within a specific UI environment.
+Furthermore, we demonstrate how this task can leverage datasets designed for the inverse problem of UI automation, utilizing Android and web datasets for our experiments.
+To benchmark this task, we compare the performance of humans and state-of-the-art models, specifically GPT-4 and Gemini-1.5 Pro, using our proposed metric.
+The results reveal that both Gemini and GPT underperform relative to human performance, underscoring the challenge of the proposed task and the significant room for improvement.
+This work highlights the importance of goal identification within UI trajectories, providing a foundation for further exploration and advancement in this area.
+Intent Understanding, Proactive Agents, Personalized Agents, Multimodal Interaction
+journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3717525isbn: 979-8-4007-1331-6/2025/04ccs: Human-centered computing User modelsccs: Human-centered computing User studiesccs: Computing methodologies Intelligent agentsccs: Human-centered computing Field studies
+1. Introduction
+Autonomous agents that interact with GUIs to complete tasks for users have drawn increasing interest (Hong et al., 2023; Gur et al., 2023; Yang et al., 2023a).
+These agents interpret user-provided instructions and iteratively interact with GUIs to complete the desired task.
+In this work, we propose empowering agents with improved personalization abilities, allowing them to identify the underlying goals of users from their observed activity within the GUI environment.
+Understanding the underlying goals and intents of users is a cornerstone of user modeling, enabling agents to proactively align their behavior with user preferences, habits, and long-term objectives.
+This capability marks a significant step toward achieving personalization, where agents can provide context-aware, effective assistance that meets individual user needs (Li et al., 2024).
+Consider the scenario in Figure 1 where a user books flight tickets for a vacation.
+An ideal agent would observe these actions, understand the underlying user goal, and then proactively suggest booking a hotel for the same dates, and make the dates visible in the calendar.
+Our work extends a long line of research on recognizing user goals from their observed behaviour, including intent, activity, and plan recognition.
+However, most prior work addressed settings in which the input consists of natural language user utterances, like search queries, dialog utterances, or social media posts, while the task was perceived as structured classification, selecting a category label from a predefined list.
+Our work, to the best of our knowledge, is the first to identify user goals from UI interactions, while providing a natural language description of the user goal as output.
+We refer to Section 2 for an extensive overview.
+In this paper, we first define the task of generating a natural language description of underlying user goals from observed UI trajectories, that is, from multi-modal (text and screen image) traces of user-system interactions.
+A challenging aspect of the task is its inherent ambiguity, since multiple user goals can often lead to the same UI activity.
+Second, we observe that our task can be framed as the inverse of the known UI automation task, where an agent needs to perform a sequence of UI actions given a natural language description of the user’s goal (Li et al., 2023; Wen et al., 2023).
+We further introduce manual and automatic evaluation protocols, assessing whether the predicted and gold task descriptions are paraphrases within the given UI context.
+Subsequently, we conducted experiments over both web and Android UI sessions, leveraging existing UI automation datasets while swapping the input and output roles.
+Over these data, we compared and analyzed the performance of humans and state-of-the-art multi-modal models, showing that there is substantial room for modeling improvements in future work.
+Overall, we offer the following contributions: (1) introducing and formalizing the task of goal identification from UI trajectories, highlighting its significance and potential applications; (2) demonstrating how existing datasets for UI automation can be adapted for this task by interpreting it as the ”reversed” process of the other; (3) proposing both manual and automatic evaluation methodologies for performance assessment; and (4) evaluating the performance of both humans and state-of-the-art models on this task.
+We propose that these contributions would trigger research on this timely challenge.
+Figure 1.
+An example of a user performing a flight booking task.
 The agent first observes the UI interactions, comprehends the task’s essence, and then offers help with related tasks, like booking a hotel and blocking calendar dates.
 We focus on the first part, comprehending the task by observing the UI interactions.
-Identifying user goals from UI interactions is challenging.
-Although ensuring a correct goal is achievable through the observed actions is essential, this alone is not enough.
-The fact that multiple tasks can often lead to the same actions making the task ambiguous, further highlighting the complexity of the task and the need for robust methods.
-Additionally, the involvement of multiple modalities, such as visual and text, increases complexity.
-In this paper, we explore the task of goal identification from observed UI trajectories.
-In particular, we introduce a robust manual and automatic evaluation protocol.
-Our proposed metric primarily assesses whether two task descriptions are paraphrases within the UI context while also enabling us to quantify the ambiguity.
-We conducted extensive experiments in both web and Android UIs, leveraging existing UI automation datasets by swapping the roles of input and output.
-We compared the performance of humans and state-of-the-art multi-modal models on this task and found that the models significantly underperform, indicating substantial room for improvement.
-Common errors and potential mitigations are highlighted, paving the way for future research.
-Overall our main contributions are as follows: (1) presenting the first study on the task of goal identification from UI trajectories, including task formulation and definitions; (2) introducing a novel manual and automatic evaluation methodology and (3) assessing the performance of both humans and models on this task, providing a comprehensive analysis of our findings.
-2 Task Definition
-Given an observed UI trajectory performed by a user with the intention to complete a certain task, our goal is to recover the user’s original intent from the observable trajectory.
-As mentioned in § 1, this intent identification setting is effectively the inverse problem of the known UI Automation task.
-We therefore adopt their input and output definitions, swapping their roles, which enables the use of UI automation datasets for the intent identification task as well.
-Accordingly, our input is a UI trajectory - a sequence of steps performed by the user.
-Each step consists of a snapshot of the UI environment at that moment, along with the corresponding action the user took at that step.
+[…]
+The figure illustrates a sequence of UI interactions in a flight booking scenario.
+A user selects flight details like destination and dates, and the agent interprets these actions.
+Additional UI interactions for related tasks, such as booking a hotel and calendar updates, are shown.
+The agent observes and understands these UI trajectories.
+2. Background and Related Work
+In this section, we provide an overview of key related research areas that aim to learn personal user behaviors and preferences, specifically intent recognition, activity recognition, and plan recognition.
+Intent Recognition
+Intent recognition, also referred to as intent classification, is a well-established field focused on identifying user intentions based on specific inputs.
+Traditionally, most research in this domain has concentrated on textual inputs, such as individual messages or utterances, which are then categorized into predefined intent classes.
+For instance, studies like (Schuurmans and Frasincar, 2019; Kuchlous and Kadaba, 2020; Larson and Leach, 2022) have worked on improving intent classification within dialogue systems across various domains, focusing on the classification of short texts or single utterances into predetermined intent categories (e.g.
+“find a train from Barcelona to Madrid” would be classified into a system intent called ”find_train”).
+Others, classified social media posts to determine whether they express an intention to make a purchase (Gupta et al., 2014; Haque et al., 2019).
+To enhance understanding of user queries and infer its underlying intent, some works leverage external knowledge sources to get better results (Hu et al., 2009).
+In addition to text-based inputs, multimodal methods have been developed to incorporate various types of inputs.
+For example, (Zhang et al., 2021) and (Gonzaga et al., 2021) investigate the use of both images and text: the former analyzes social news content to identify marketing intents and to classify intent topics, while the latter combines image and textual data to classify social media posts, aiming to identify the writers’ intent such as provocative, informative, promotive and more.
+Zhang et al.
+(2022) introduces a multimodal dataset for intent recognition in TV series, where inputs include visual, auditory, and textual data, while the outputs are classified into one of 25 possible intent categories.
+Activity Recognition
+Activity recognition, as defined by Sukthankar et al.
+(2014) and closely related to intent recognition, involves identifying specific human activities based on a series of observations, often through sensor data.
+While intent recognition focuses on understanding user intentions, activity recognition mostly concerned with classifying physical actions.
+Typically, systems in this domain are developed to classify these activities using sensory inputs.
+Khan et al.
+(2022) uses a neural network trained on 2D skeletal data captured by a motion sensor to classify human poses; Ijaz et al.
+(2022) integrated accelerometer signals along with skeletal data to recognize and categorize nurse activities into 12 distinct types.
+A prominent study (Kwapisz et al., 2011) developed a supervised learning algorithm that uses accelerometer signal from Android smartphones to classify physical activities like walking, jogging, sitting, and standing.
+Moreover, research such as (Liao et al., 2005) utilizes GPS traces to classify human activity and label significant locations, employing relational Markov networks to achieve this.
+Plan Recognition
+In a plan recognition problem, a system is given a series of actions performed by an agent and is expected to infer the overall plan or goal which explains those actions (Kautz et al., 1991).
+The distinction between activity recognition and plan recognition is the difference between recognizing a single activity and recognizing the relationships between a set of such activities that result in a complete plan or goal.
+In this field, foundational works like (Charniak and Goldman, 1993; Goldman et al., 2013) rely on structured inputs and a plan library — a collection of possible plans and their associated goals — to construct probabilistic models that infer the plan from observed actions.
+Alternative approaches, such as in (Hong, 2001), develop algorithms that do not depend on plan libraries; instead, they utilizing in-domain knowledge and applying this to UNIX system commands to understand broader user goal.
+A hybrid approach is demonstrated (Granada et al., 2017), where the authors combine CNN models to detect human activities from a video stream of meal preparation and apply symbolic reasoning with a plan library to recognize the overall plan (e.g., actions like ”breaking eggs” and ”mixing ingredients” might be recognized as part of a plan to create an omelet).
+3. Task Definition
+Given an observed UI trajectory (see below) performed by a user with the intention to complete a certain task, our goal is to recover the user’s original intent from the observable trajectory.
+As mentioned in § 1, this setting is effectively the inverse problem of the known UI Automation task.
+We therefore adopt their input and output definitions, swapping their roles, which enables the use of UI automation datasets for our task as well.
+Accordingly, our input is a UI trajectory — a sequence of individual interaction steps between the user and the system along the session.
+Each step consists of a snapshot of the UI content at that moment, along with the corresponding action the user took at that step (see Appendix A for the specific trajectory formats).
 From this trajectory, our goal is to generate a natural language description that accurately captures the user’s intended task.
 Within the scope of this paper, we address the core setting of the intent identification task and therefore assume that the observed UI trajectory indeed successfully fulfills the underlying user intent.
-The intent identification task, similar to other text generation tasks like summarization where multiple valid outputs can exist, is inherently ambiguous, mostly because the same trajectory may fulfill multiple intents.
+The intent identification task, similar to other text generation tasks like summarization where multiple valid outputs can exist, is inherently ambiguous, mostly because the same trajectory may fulfill multiple intents, often due to varying specificity levels of the original user intent (as captured in the dataset).
 For example, when observing a Sushi restaurant booking, the user might have asked for that specific Sushi place, or more broadly for some restaurant in that area, or of that cuisine type.
-When a model identifies intent from trajectory, we expect it to predict the most likely one.
-3 Evaluation Methodology
+When a model identifies an intent from the given trajectory, we expect it to predict the most likely one, as reflected in the dataset distribution.
+4. Evaluation Methodology
 This section outlines our evaluation methodology.
 Given an input UI trajectory and a corresponding gold task description, we assess whether a predicted task description matches the gold reference.
-We start with necessary definitions (Section 3.1), followed by a human evaluation protocol (Section 3.2) and a language model-based automatic evaluation (Section 3.3).
-3.1 Definitions
+We start with necessary definitions (Section 4.1), followed by a human and automatic evaluation protocols (Sections 4.2 and 4.3).
+4.1. Definitions
 Task Fulfillment by a Trajectory
@@
 (2023), we differentiate between fulfillment for information-seeking intents and transactional intents (e.g.
-purchasing an item, booking a flight or changing settings).
+purchasing an item or changing settings).
 For the latter, fulfillment is achieved upon successfully completing the specific requirement outlined in the task.
 For information-seeking intents, fulfillment is achieved when the trajectory provides the necessary information sought in the user intent.
-We note that a fulfilling trajectory may provide some additional information beyond the intent, if such additional information is inherently bundled in the UI environment together with the sought information (e.g.
-some UI environments may display review scores along with other related information).
+We note that a fulfilling trajectory may provide some extra information beyond the intent, if such additional information is inherently bundled in the UI environment together with the sought information.
 Satisfaction Relation between Tasks
 We next aim to specify a matching criterion between two task descriptions, specifically a predicted one and the corresponding gold-reference.
-Given two task descriptions A and B and a UI environment, we say that A satisfies B in that environment if every reasonable user trajectory that fulfills A would also fulfill B.
+Given two task descriptions A and B and a UI environment, we say that A satisfies B in that environment if every reasonable trajectory that fulfills A would also fulfill B.
 In essence, this means that completing task A necessarily results in completing task B, making B a more general task than A in that UI environment.
-For instance, the task “Purchase the earliest train ticket to Edinburgh” satisfies the task “Purchase a ticket to Edinburgh” but not vice versa.
+For instance, the task “Purchase the earliest train ticket to Edinburgh” satisfies the task “Purchase a train ticket to Edinburgh” but not vice versa.
 Building on these definitions, we consider a predicted task description to successfully match the gold description if the two mutually satisfy each other, and partial match when only one satisfies the other.
@@
 We highlight the relation between match and ambiguous trajectory, which fulfills multiple task descriptions that do not match each other.
-3.2 Human Evaluation Protocol
+4.2. Human Evaluation Protocol
 As with many text generation tasks, human evaluation is essential due to the limited reliability of automatic evaluation metrics.
-In our case, the human evaluator observes the gold and predicted task descriptions, along with the corresponding trajectory, and assesses (1) whether the trajectory fulfills the predicted task, and (2) whether each of the predicted and gold task descriptions satisfies the other.
-To verify the quality of our suggested metric, we sampled randomly 50 instances from Mind2Web dataset (see § 4.1) and generated task descriptions using two baseline models (see § 4.2) evaluated in our experiments.
-We measured the pairwise inter-annotator agreement between three of the authors yielding average Cohen’s Kappa of 0.79 for fulfillment judgement and 0.77 for satisfaction judgement.
-These agreement levels are in the higher range of the “Substantial agreement” level, thus we proceeded to manually evaluate each baseline model in our experiments § 5 by a single annotator.
-3.3 Automatic Evaluation Metric
+In our case, the annotator observes the gold and predicted task descriptions, with the corresponding trajectory, and assesses (1) whether the trajectory fulfills the predicted task, and (2) whether each of the predicted and gold task descriptions satisfies the other.
+Additionally, annotators validated the given data instance, excluding from the evaluation noisy instances in the original datasets (details in Appendix B).
+To assess the quality of our proposed metric, we randomly sampled 50 instances from each of our datasets (see § 5.1) and generated task descriptions using our two baseline models (see § 5.2).
+We then measured the pairwise inter-annotator agreement among three of the authors, resulting in an average Cohen’s Kappa of 0.79 and 0.77 for fulfillment and satisfaction judgments, respectively, in the web dataset, and 0.91 and 0.86 in the Android dataset.
+These agreement levels are considered high according to Kappa values, which justified our decision to manually evaluate each baseline model in our experiments using a single annotator.
+4.3. Automatic Evaluation Metric
 We propose utilizing a Large Multimodal Model (LMM) as an automatic evaluator for the satisfaction criteria.
-Recent advancements in LMMs, such as those demonstrated in He et al.
-(2024) and Pan et al.
-(2024), have shown promising results in employing GPT-4 Achiam et al.
-(2023) to assess task completion by autonomous agents.
-Building on this, we leverage the latest GPT-4o model as the automatic evaluator to determine whether two task descriptions, are mutually satisfied in the context of the trajectory.
-The specific prompt used is detailed in Appendix D.
-Our analysis indicates that the proposed automatic evaluation generally agrees with the human annotators, with an F1 score of 0.75, and can be a sufficient substitute for manual evaluation, enabling quick development cycles.
-Dataset Non Partial  Match ($\uparrow$)
- Match  Match
+Recent advancements in LMMs, such as those demonstrated in (He et al., 2024) and (Pan et al., 2024), have shown promising results in employing GPT-4 (Achiam et al., 2023) to assess task completion by autonomous agents.
+Building on this, we leverage the latest GPT-4o model as the automatic evaluator, to determine whether two task descriptions are mutually satisfied in the context of the trajectory (prompt details are in Appendix D).
+Measuring agreement with our human evaluation yielded a Kappa value of 0.48 (moderate agreement), suggesting a potential utility of model-based automatic evaluation for development cycles while highlighting the need for manual evaluation.
+Table 1.
+Manual evaluation of human-generated task descriptions, at the different levels of match criteria (determined by the satisfaction relation) between annotators (Hi) and gold task descriptions (G), as well as among annotators.
+(H) represents average scores.
+Dataset Non Match Partial Match Match (↑)
 Mind2Web H vs. G 0.07 0.13 0.80
    H1 vs. H2 0.09 0.10 0.81
-AitW H   vs. G 0.12 0.32 0.56
-  H1 vs. H2 0.05 0.20 0.75
-Table 1: Manual evaluation of human generated task descriptions, at the different levels of match criteria (determined by the satisfaction relation) between annotators (Hi) and gold task descriptions (G), as well as among annotators.
-(H) represents average scores.
-4 Datasets and Baseline Models
+AitZ H vs. G 0.02 0.22 0.76
+   H1 vs. H2 0.02 0.04 0.94
+5. Datasets and Baseline Models
+Table 2.
+Manual and automatic evaluation scores of fulfillment relation and the different levels of match criteria (determined by the satisfaction relation), between model predictions and gold task descriptions.
 Manual Eval Automatic Eval
-Dataset Model Fulfillment Non Match Partial Match  Match ($\uparrow$) Non Match Partial Match  Match ($\uparrow$)
+Dataset Model Fulfillment Non Match Partial Match Match (↑) Non Match Partial Match Match (↑)
 Mind2Web GPT-4 0.86 0.14 0.42 0.44 0.18 0.40 0.42
 Gemini-1.5 0.87 0.20 0.22 0.58 0.18 0.28 0.54
-AitW GPT-4 0.65 0.50 0.20 0.30 0.62 0.19 0.19
-Gemini-1.5 0.89 0.19 0.31 0.50 0.36 0.29 0.35
-Table 2: Manual and automatic evaluation scores of fulfillment relation and the different levels of match criteria (determined by the satisfaction relation), between model predictions and gold task descriptions.
-4.1 Evaluated Datasets
-Given that our task is the inverse of UI automation task, leveraging the datasets created for UI automation is a natural choice.
-In these datasets, humans perform actions within a UI environment, based on a given task description.
-For our experiments, we explore two UI environments: web and Android.
-We utilize the Mind2Web dataset Deng et al.
-(2023) for the web environment, as it is the most widely used benchmark for autonomous web agents.
+AitZ GPT-4 0.78 0.22 0.19 0.59 0.32 0.34 0.34
+Gemini-1.5 0.88 0.17 0.26 0.57 0.37 0.27 0.36
+5.1. Evaluated Datasets
+A suitable dataset for our task should contain natural language descriptions of user intents, along with traces of UI sessions that fulfill them.
+Aiming to satisfy this requirement, leveraging datasets that were created originally for UI automation is a natural choice.
+In these datasets, each instance was created by first specifying a particular user intent, followed by performing and recording a UI session that fulfills the given intent.
+Thus, a suitable dataset for our task of intent recognition is obtained from a UI automation dataset by swapping the input and output roles.
+To that end, we explored several existing UI automation datasets but found that many of them do not fully align with the requirements of generic real-world intent recognition.
+For example, WebGUM operates in a simplified GUI environment while WebShop (Furuta et al., 2023; Yao et al., 2022) is limited to the shopping domain.
+RUSS and WebArena (Xu et al., 2021; Zhou et al., 2023), while providing user intents or instructions, do not possess the needed UI traces.
+PixelHelp and OmniACT (Li et al., 2020; Kapoor et al., 2025) provide only UI-level procedural instructions rather than natural high-level intents which are independent of a particular UI environment (asking to click a button, filling a text field etc.).
+Lastly, RICO (Deka et al., 2017) lacks annotations for user intents.
+Considering these limitations, we turned to datasets that do satisfy our task’s requirements, focusing on two UI environments: web and Android.
+For the web environment, we utilize the Mind2Web dataset (Deng et al., 2023), the most widely used benchmark for autonomous web agents.
 The dataset was created by curating diverse tasks across popular websites, with annotators performing a series of actions to complete the goal.
-For Android, we used the Android in the Wild (AitW) Rawles et al.
-(2023) dataset, notable for its diverse task descriptions spanning various domains.
-Further details in Appendix A.
-4.2 Models
+For Android, we used the prominent Android in the Wild (AitW) (Rawles et al., 2023) dataset, while focusing on its quality-filtered subset Android in the Zoo (AitZ) (Zhang et al., 2024) (details in Appendix A).
+5.2. Models
 Given the multimodal nature of UI trajectories, our models must be adept at interpreting both text and images.
-We selected two state-of-the-art LMMs, Gemini 1.5 Pro Reid et al.
-(2024) and GPT-4 Achiam et al.
-(2023).
-These models are at the forefront of handling combined text and image inputs and offer a sufficiently large context window, which is necessary for our experiments.
-Our experiments focused on in-context learning, employing both zero-shot and few-shot prompting.
-The model was guided through a Chain-of-Thought Wei et al.
-(2022) process, to analyze the trajectory in a step-by-step manner (details in Appendix D).
-In web, inspired by SeeAct Zheng et al.
-(2024) we drew a red bounding box around the element that the user interacted with to guide the model’s attention, as well as to break ambiguity where textual action descriptions were not sufficient.
-Additionally, we truncated lengthy web page images based on the bounding box position to further focus the model’s attention (details in Appendix B).
-For Android, no special modifications were necessary, as actions were overlaid on the screenshots and the screenshots were limited to the visible area.
-5 Experiments
-5.1 Human Performance Evaluation
-To asses task difficulty and establish a baseline for models, two NLP practitioners, unfamiliar with the datasets, independently composed 50 task descriptions from Mind2Web, and 50 from AitW.
-An annotator then evaluated [^1] During the evaluation it was observed that human-generated tasks were inherently fulfilled by the trajectory.
-Therefore, we do not report this score.
-these descriptions as explained in § 3.2.
-The results, summarized in Table 1, reveal more matches between the human annotators and the gold task descriptions in Mind2Web compared to AitW.
-Upon analysis of the AitW disagreements, we found that the gold descriptions were often more specific than those provided by human annotators.
-In some cases, the gold tasks were already fulfilled by the trajectory, resulting in no clear interactions that indicate the user goal.
+We selected two state-of-the-art LMMs, Gemini 1.5 Pro (Reid et al., 2024) and GPT-4.
+These models are at the forefront of handling combined text and image inputs and offer a sufficiently large context window for our experiments.
+In our experiments, the model was guided through a Chain-of-Thought (Wei et al., 2022) process, to analyze the trajectory in a step-by-step manner (prompt in Appendix D).
+In web, inspired by SeeAct (Zheng et al., 2024) we drew a red bounding box around the element that the user interacted with to guide the model’s attention, as well as to break ambiguity where textual action descriptions were not sufficient.
+To further focus models, we also truncated lengthy web images based on the bounding box position.
+(details in Appendix B).
+For Android, no special modifications needed, actions are overlaid on the screenshots, and the screenshots are naturally smaller.
+6. Experiments
+6.1. Human Performance Evaluation
+To asses task difficulty and establish a baseline for models, two NLP practitioners, unfamiliar with the datasets, independently composed 50 task descriptions from Mind2Web, and 50 from AitZ.
+An annotator then evaluated [^1] During the evaluation it was observed that human tasks were inherently fulfilled by the trajectory.
+them as explained in § 4.2.
+The results, summarized in Table 1, reveal more matches between the human annotators and the gold task descriptions in Mind2Web compared to AitZ.
+Upon analysis of the disagreements, we found that the gold descriptions in Android were often more specific than those provided by human annotators.
+In most cases, the gold tasks were already fulfilled by the trajectory, resulting in no clear interactions that indicate the user goal.
 For example, if the task was “Turn WiFi on” and the WiFi was already on, the annotators inferred a more general task, such as “Show WiFi settings”.
-Further analysis of non-matching records across both datasets revealed that ambiguous task descriptions contributed to disagreements between raters.
-Detailed analysis and examples can be found in Appendix C.
-5.2 Model Performance Evaluation
-Model evaluation included manual assessment of 100 predicted task descriptions and automated evaluation of 1,000 predictions per model and dataset (details provided in Appendix B).
-The results, outlined in Table 2, reveal that Gemini outperforms GPT-4 on both datasets, achieving higher match scores but still falling short of human performance.
-In AitW, GPT-4 often generated too abstract goals such as “Explore the main page of a shopping app” resulting in higher non-match scores.
-In contrast, Gemini tended to be more specific and heavily influenced by the content on the screen.
+Further analysis of non-matching records across both datasets revealed that ambiguous trajectories caused human disagreements.
+Detailed analysis in Appendix C.
+6.2. Model Performance Evaluation
+Model evaluation included manual assessment of 100 predictions from each dataset and automatic evaluation using the entire test sets: 1,013 from Mind2Web and 506 from AitZ (see Appendix B).
+The results, outlined in Table 2, reveal that Gemini outperforms GPT on the Mind2Web dataset, achieving higher match scores but still falling short of humans.
+In AitZ, both models perform comparably.
+Gemini tends to be more specific and detailed than the gold references, while GPT often generates more general, abstract goals.
 Both models, in certain instances, misidentified the actual user intent (e.g.
@@
 “Watch The Dark Knight trailer”) and exhibit limitations in visual screen understanding, leading to missing details, incorporating irrelevant information and hallucinations of non-existent information.
-We provide a detailed error analysis and further discussion in Appendix C.
-Experiments with few-shot prompting did not yield improved results.
-We hypothesize that the large context size, resulting from including multiple trajectories with several images each in the prompt, were too complex models to leverage.
-6 Conclusion and Future Work
-In this paper, we introduced a novel task for UI-related agents by identifying user goals from UI trajectories and proposed a reliable evaluation methodology.
-Our experiments on Android and web datasets reveal a significant gap between humans and state-of-the-art models, with a detailed analysis provided for this discrepancy.
+Detailed discussion in Appendix C.
+These results underscore the complexity of accurately capturing user intent.
+Our experiments also concluded few-shot learning.
+However, in our setting, this is challenging due to multiple images, textual inputs and the thought process within example.
+Adding a single example deteriorated results, likely because the large context size and the difference between the exemplar and the test trajectory.
+7. Conclusion and Future Work
+In this paper, we introduced the novel task of identifying user goals from UI trajectories and proposed a reliable evaluation methodology.
+Our experiments on Android and web datasets reveal a significant gap between humans and state-of-the-art multi-modal models, which is analyzed in detail.
 Future work may include fine-tuning experiments and enhancing visual understanding for UI environments.
-We propose testing our task by evaluating how well it improves downstream tasks like agent personalization and completions.
-Additionally, expanding the scope to other GUIs like iOS and Windows would broaden the applicability and impact of our findings.
+We propose testing our task by evaluating its utility on downstream tasks like agent personalization and suggestions.
+Additionally, expanding the scope to other GUIs like iOS and Windows would broaden the impact of our findings.
 Limitations
@@
 It is essential for researchers and developers to proactively address these concerns to foster trust and responsible innovation in the field of autonomous agents.
-Acknowledgements
+Acknowledgements.
 We wish to thank Guy Rom, Xiang Deng, and Eliya Nachmani for their valuable feedback on our research.
@@
 Appendix A Datasets Overview
-Figure 2: An instance from Mind2Web, representing a full trajectory accomplishing the task description above.
-Figure 3: An instance from AitW, representing a full trajectory accomplishing the task "Set an alarm". The blue plus sign indicates the area on the screen where the tap occured.
-Mind2Web
+Figure 2. An instance from Mind2Web, representing a full trajectory accomplishing the task description above.
+[…]
+The figure depicts a sequence of user interactions within a web environment to complete the task of ”Find unique experiences in London of maximum one hour duration which are rated four stars and above”.
+It consists of six sub-images, each showcasing an web screen representing a specific step in the user’s actions.
+This example is taken from the Mind2Web dataset.
+Figure 3. An instance from AitW, representing a full trajectory accomplishing the task ”Set an alarm”. The blue plus sign indicates the area on the screen where the tap occurred.
+[…]
+The figure depicts a sequence of user interactions within an Android environment to complete the task of ”Set an Alarm.” It consists of five sub-images, each showcasing an Android screen representing a specific step in the user’s actions.
+This example is taken from the Android In the Wild dataset.
+A.1. Web
 We aim to give a brief overview of how Mind2Web was collected, and its format in more depth.
@@
 Figure 2 demonstrates a single instance of the data, where textual action descriptions are presented below the corresponding screenshot.
-Android-In-The-Wild
+A.2. Android
 The Android in the Wild (AitW) dataset stands out due to its extensive variety of tasks, covering 4 domains: Google-Apps, Install, Web-Shopping, and General, as well as a single-step domain excluded from this paper’s analysis.
 The dataset consists of a substantial collection of high-level instructions, trajectories of varying lengths, and a notable variety of apps and websites.
-A key strength of the AitW dataset lies in the diversity of task instructions, which provided in-depth insights into the complexities of understanding user goals across different domains.
-The high-level goal instructions were sourced from various sources, including humans (both crowd-sourced and the authors themselves), LLM-generated prompts, and technical documentation such as PixelHelp Li et al.
-(2020).
+The Android in the Zoo (AitZ) dataset was sampled from AitW to reduce tasks redundancy and to filter erroneous data points, resulting in more unique and higher-quality task descriptions.
+The high-level goal instructions in AitW were sourced from various sources, including humans (both crowd-sourced and the authors themselves), LLM-generated prompts, and technical documentation such as PixelHelp (Li et al., 2020).
 The creation process involved human annotators performing tasks on Android emulators, with their gestures being recorded.
-The AitW dataset contains 700k+ episodes comprising 30k unique instructions.
-These episodes were recorded on mobile devices running four different versions of Android.
+The episodes were recorded on mobile devices running four different versions of Android.
 Each episode contains natural language instructions and observation-action pairs.
 The observations are screenshots, while the actions are one of three types: tap, drag, or typing.
-Gesture actions are represented as taps and drags at specific <x,y> coordinates on the screen.
+Gesture actions are represented as taps and drags at specific ¡x,y¿ coordinates on the screen.
 In our utilization of this data, we presented the models with a series of screenshots with the actions drawn on them, as shown in Figure 3.
 Appendix B Experiments
-Models Configuration
+B.1. Models Configuration
 For task description prediction, we used the most recent version of Gemini, namely Gemini 1.5 Pro, updated in Vertex AI [^2] https://cloud.google.com/vertex-ai as of May 2024, with a sampling temperature of 1.0, and GPT-4-Turbo [^3] https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 (version gpt-4-turbo-2024-04-09) with a sampling temperature of 0.6.
 For automatic evaluation, we utilized the latest release from the GPT family, known as GPT-4o.
-Data
-For manual evaluation, we sampled 100 random data points from the test sets of both AitW and Mind2Web.
-For automatic evaluation, we used the entire Mind2Web test set, consisting of 1013 examples, and randomly sampled 250 examples from each of the four domains in AitW, resulting in a total of 1000 examples (as described in Appendix A).
-Web
+B.2. Data
+For the manual evaluation, we randomly sampled over 100 data points from the both Android and Web datasets.
+We then prompted models to predict the user goal.
+During the evaluation, we conducted a thorough verification process to ensure that both the gold references and the trajectories were of high quality.
+We rejected instances where the trajectory did not fulfill the gold reference, which occurred more frequently for Android.
+Additionally, for web-based tasks, we excluded cases where the screenshot was not rendered properly, making it difficult to interpret the user’s action, as well as instances where the bounding box was empty.
+As a result of this verification process, 17% of the Android examples and about 5% of the web examples were rejected.
+Ultimately, as a result of this process, the samples that were manually evaluated in Sections 4.2 and 6 were of high quality and free from such issues.
+For the automatic evaluation, as described in 6, we utilized the entire test sets from both datasets.
+B.3. Web
 We encountered a technical challenge with the Mind2Web data due to the nature of its image captures.
@@
 This was achieved by utilizing the bounding box metadata provided by Mind2Web.
-Similar to the approach taken by Zheng et al.
-(2024) and introduced by Yang et al.
-(2023a), we drew the red bounding box around the interaction element to guide the model’s attention.
+Similar to the approach taken by (Zheng et al., 2024) and introduced by (Yang et al., 2023b), we drew the red bounding box around the interaction element to guide the model’s attention.
 Additionally, adding the bounding box helped resolve ambiguities at times that the textual action description is not sufficient.
@@
 While in other cases, action descriptions exist, but the textual description matches multiple element descriptions and thus results in ambiguity that is only resolved by drawing the bounding box.
-For example, a button labeled "Add to Cart" is typically associated with each item in a web shopping list.
+For example, a button labeled ”Add to Cart” is typically associated with each item in a web shopping list.
 Without the bounding box, it is impossible to determine which specific button was clicked.
-Android
+B.4. Android
 In our efforts to replicate experiments from web, we encountered challenges due to differences in data format.
@@
 AitW’s provided visualization tools to draw the actions, also included labels of special actions such as the back button, home button, and enter.
-As well as a special "status" action: either Task Complete or Task Impossible.
+As well as a special ”status” action: either Task Complete or Task Impossible.
 However, we found that the label Task Complete often confused the models.
@@
 Appendix C Error Analysis
-Figure 4: An illustration in which the user chose a specific car primarily for its 12-inch feature, but since it was also the cheapest, annotators incorrectly assumed cost was the deciding factor.
+Figure 4. An illustration in which the user chose a specific car primarily for its 12-inch feature, but since it was also the cheapest, annotators incorrectly assumed cost was the deciding factor.
+[…]
+The figure presents an example from the Mind2Web dataset, showcasing four different trucks, each with distinct wheel sizes and prices.
+The image is annotated with the user’s selection: the second truck, which has the lowest price and features 12” wheels.
 In this section, we aim to provide a more detailed error analysis with respect to model and humans.
-Web
+C.1. Web
 In our error analysis of 30 mismatched task descriptions, distinct patterns emerged between GPT and Gemini models.
@@
 Although not exhaustive, this highlights that most trajectories in Mind2Web are probably non-ambiguous.
-Figure 5: Comparison of "Match" proportions between Gemini 1.5 Pro and GPT-4-Turbo models across different AitW domains
-Android
+Figure 5. Comparison of ”Match” proportions between Gemini 1.5 Pro and GPT-4-Turbo models across the different domains
+[…]
+The figure presents a histogram comparing the ”Match” metric values across different domains in the Android in the Zoo dataset.
+The domains include ”General,” ”Google Apps,” ”Install,” and ”Web Shopping.” The comparison is made between the GPT-4 and Gemini 1.5 Pro models.
+C.2. Android
 Comparing GPT and Gemini’s performance on the Android dataset revealed notable variations across different domains.
-Gemini consistently surpassed GPT in all domains, primarily due to GPT’s tendency to offer abstract goals and frequent random hallucinations.
-Gemini’s errors varied based on the specific domain.
-Both models were proficient in the “General” domain, but faced challenges in “Web Shopping” and “Google Apps”.
+Both models were proficient in the “General” domain, but faced challenges in “Web Shopping” and “Install”.
 This disparity is due to the nature of the “General” dataset, which contains search queries that explicitly reveal the user’s intent.
@@
 Appendix D Instructions and Prompts
-Figure 6: Instructions for human-annotators to conclude if a task is fulfilled by a trajectory, and if two task descriptions satisfying each other.
+Figure 6.
+Instructions for human-annotators to conclude if a task is fulfilled by a trajectory, and if two task descriptions satisfying each other.
 These instructions were the core prompt of the automatic evaluator.
-Figure 7: Goal identification annotation task instruction to measure human performance.
-Figure 8: Instruction used to guide GPT and Gemini when predicting a task description given a UI trajectory. We swap AitW section and Mind2Web according to the input.
-Figure 9: Model instructions for evaluating Satisfaction relation between two task descriptions (A and B), given a corresponding trajectory.
+[…]
+The figure contains text that was used to instruct human participants on how to evaluate the fulfillment and satisfaction relations when comparing two task descriptions with a UI trajectory.
+Figure 7. Goal identification annotation task instruction to measure human performance.
+[…]
+The figure contains text that was used to instruct human participants while performing the task of identifying user goals from UI trajectories.
+Figure 8. Instruction used to guide GPT and Gemini when predicting a task description given a UI trajectory. We swap AitW section and Mind2Web according to the input.
+[…]
+The figure contains text that was used to guide Gemini 1.5 Pro and GPT-4 models in predicting user goals based on UI trajectories.
+Figure 9.
+Model instructions for evaluating Satisfaction relation between two task descriptions (A and B), given a corresponding trajectory.
 Includes web/mobile data format instructions and few-shot examples for experimentation (detailed on the next page).
-Figure 10: Few-Shot exemplars demonstrating the Satisfaction Relation Prompt (Figure 9). These simplified examples highlight the nuances of task satisfaction in real-world scenarios.
+[…]
+The figure contains a table with text displaying the instructions used to guide the GPT-4 model in automatically rating the satisfaction relation between two task descriptions.
+Figure 10. Few-Shot exemplars demonstrating the Satisfaction Relation Prompt (Figure 9). These simplified examples highlight the nuances of task satisfaction in real-world scenarios.
+[…]
+The figure contains a table with text, displaying few-shot examples used in the prompt to guide the GPT-4 model in automatically rating two task descriptions.